OpenStudy (anonymous):

Solve for the variable. x + 80 = -19 A. 99 B. -99 C. -61 D. 61

undeadknight26 (undeadknight26):

80 - (-19) = ???

OpenStudy (adilalvi):

b

undeadknight26 (undeadknight26):

no -_-

OpenStudy (twopointinfinity):

-19 +80 = x

OpenStudy (anonymous):

@undeadknight26 99 right

OpenStudy (twopointinfinity):

80 - 19 = x

OpenStudy (adilalvi):

hows not lol its so simple just shift 80 on right side it will become -80 then add with -19.

OpenStudy (anonymous):

GJ UNDEAD!!

OpenStudy (anonymous):

brei x+80=-19 if you subtract 80 from both sides you are left with -99.

undeadknight26 (undeadknight26):

i believe so... 80 - (-19) = 80 + 19 = 99

OpenStudy (anonymous):

No when you shift numbers over, it takes the opposite sign, so undeadknight26 that's not right

OpenStudy (anonymous):

undead your figorgetting that you have to get the x by itself. what your doing would leave you with x+99=0

OpenStudy (anonymous):

@mrrodriguez -99?

OpenStudy (anonymous):

therefore no matter if you do it like you said @undeadknight26 you still have to get the x by itself so you would have to shift the 99 over which you do by subtracting it. leaving you with -99. all you did was as an extra step

OpenStudy (anonymous):

That is correct

undeadknight26 (undeadknight26):

aye...i was wrong...thanks for reminding me that guys :D

OpenStudy (anonymous):

@undeadknight26 it's the thought that counts

undeadknight26 (undeadknight26):

thanks miss brei!

OpenStudy (anonymous):

that is true. lol but hey you were half right @undeadknight26 just left one step out lol